OpenStudy (anonymous):

MEDAL simplify x^2+2x-35/x^2+4x-21

OpenStudy (anonymous):

@neer2890 are you abkle to help with this one?

OpenStudy (neer2890):

you have to use factorization method for this one for the equation ax^2+bx+c=0 we always factorize the acx^2 such that it becomes bx for e.g., x^2+2x-35=0 x^2+7x-5x-35 =x(x+7)-5(x+7) =(x+7)(x-5)

OpenStudy (neer2890):

here when we multiply x^2 and -35 its -35x^2 also when we multiply 7x and -5x it becomes -35x^2

OpenStudy (neer2890):

using same method for x^2+4x-21 this becomes (x+7)(x-3) so when we divide both of them (x+7) cancels out and it remains \[\frac{ (x-5) }{ (x-3) }\]

OpenStudy (tkhunny):

You have written x^2+2x-35/x^2+4x-21 = \(x^2 + 2x - \dfrac{35}{x^{2}} + 4x - 21\) Is that what you meant? If not, you may wish to add some parentheses to make it say what you want.

OpenStudy (neer2890):

alternatively you can use this formula for finding the roots of the quadratic equation \[ax ^{2}+bx+c=0\] \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]